# COSMIC: Mutual Information for Task-Agnostic Summarization Evaluation

### v2

@@ 5.1 Downstream tasks
Classification tasks. We selected four different classification tasks and corresponding classifiers to be run on the source texts and summaries and compared their outputs. We report the probability of error (6), i.e., the classifier outputting a different label on the source text and the summary. We report the results of three tasks: sentiment analysis, policy classification, Emotion classification and ChatGPT detector performed with open-source models from the HuggingFace hub.
@@ 5.2 Correlation with Learnt metrics
There are not many available metrics trained on human judgement for summarization. We chose to evaluate our metric on the SEAHORSE metrics Clark et al. (2023) as they are the most recent and provide interpretable metrics.
Seahorse metrics. These metrics — learnt from human judgement — assess summaries along $6$ axes: Main ideas; Attribution; Comprehensible; Grammar; Repetition and Conciseness. Main ideas and Attribution, respectively, measure if the main ideas of the source text are present in the summary and if the information in the summary does indeed come from the source text. Comprehensible and Grammar measure if the summary is fluent and grammatically correct, while Repetition and conciseness measure, if there are NO repetitions and the summary, is concise. The $\Pr(\text{Yes})$ metric corresponds to the average probability over the dataset that the SEAHORSE model predicts the answer Yes to the corresponding question. While we would not expect our MI metric to correlate with the Grammar or Comprehensible scores, it should strongly correlate with the Main Ideas and Attribution scores as these are proxies to the information contained in the summary.
@@ 6 Experimental Results
Table 1: Established (reference-free) quality metrics Spearman correlation with performance on the downstream classification tasks. Where Sent. analysis stands for sentiment analysis, GPT det. for GPT detector, Policy for policy classification, Emotion for emotion classification and Emb. for paraphrase embedding.
@@
$I(\mathbf{S};\mathbf{T})$ 0.65 0.73 0.83 0.72 0.79
Attribution 0.68 0.83 0.40 0.68 0.76
Main ideas 0.34 0.39 0.69 0.40 0.60
BARTScore 0.68 0.86 0.46 0.72 0.74
BERTScore 0.70 0.89 0.77 0.75 0.69
ROUGE-L 0.54 0.65 0.71 0.58 0.29
@@
Correlation with downstream tasks performance. In Table 1, we report the correlation between the different metrics and the error probability for different classification tasks and with the dot product for the Paraphrase embedding task. The MI is competitive with both the common quality estimation metrics such as BERTScore and BARTScore. In addition, we report the correlation of the metrics trained on human judgement from the SEAHORSE benchmark to predict whether the main idea of the text is present in the summary and if all elements of the summary are attributable to the source text.
@@
Figure 2: Spearman correlation with human judgement estimated by the SEAHORSE metrics. As one would expect the MI does correlate with Attribution and Main ideas but not with comprehensible, grammar or repetition.
@@
Correlation with human-judgement-based metrics. As the SEAHORSE metrics are trained to mimic human judgement, we can use them to assess the behaviour of the MI. We found consistent and expected correlations with the different SEAHORSE metrics (Figure 2). The mutual information correlates well with Attribution and Main Ideas but not with Comprehensible, Grammar and Repetition. This is not surprising as one would expect the mutual information to capture the amount of information in the summary that is attributable to the source text and not the grammatical correctness or fluency of the summary. A rather surprising result is the high correlation with Conciseness. We believe this is because the conciseness correlates with the strength of the summarizer, which correlates with the MI between the source texts and the summaries. The stronger the language model, the better the encoding of the source texts as a summary is. It is also plausible that the learnt metrics are flawed in some ways hindering the results.
@@
We also observe that the MI correlates more consistently with these Main idea, Attribution and Conciseness scores than the common quality assessment metrics, e.g. BERTScore and BARTScore (see Figure 2 for further details). It suggests that in addition to being a good predictor of the performance of the downstream tasks, the MI is also a better predictor of the human judgement of the quality of the summaries. It is notable that the MI, a theoretical quantity derived from Shannon’s MI reproduces human judgement expectations without any training on human judgement data.
@@
Comparing summarizers with COSMIC. In Figure 3, we compare the MI of different models and report their size. We observe that OOD models — trained on Arxiv or medical data — perform very poorly whereas IN distribution models such as BART display significantly higher MI. Interestingly enough, the size of the model does not seem to be a good predictor of the MI. In Appendix F we explore further use of the mutual information to compare the informativeness of different summarizers between themselves to construct a hierarchy based on their mutual informativeness.
@@
Figure 3: Comparison of models of different sizes and strengths, finetuned on different datasets regarding the MI between their summaries and the source text on different evaluation datasets. The dotted lines highlight the highest MI for each dataset reached by a model and the grey area represents the number of parameters of the model
7 Discussion of Numerical Results
@@
Our findings reveal that various conventional metrics do not consistently align with the effectiveness of downstream tasks. As illustrated in Figure 4, different metrics exhibit distinct behaviors and correlations with other metrics and the performance of downstream tasks.
@@
BERTScore. BERTScore does not correlate consistently with human-judgement-based metrics which is consistent (Figure 2) but displays nonetheless a good correlation with the performance of the downstream tasks. It suggests that the MI is not only better grounded theoretically but also an all-around more consistent metric.
Figure 4: Correlation between the main metrics and different performance metrics for the different datasets. The MI (green) closely follows the behaviour of the Attribution metric but is a better predictor of the performance of the downstream tasks (Sentiment analysis, GPT Detector, Topic classification, etc). By contrast, ROUGE scores do not display consistent correlations. For metrics to be considered effective, they should consistently exhibit positive correlations with each other. We report in Section E.3 the aggregated numerical results.
@@
Behaviour of MI. The MI displays a very similar behaviour to the metric trained to detect whether the ideas presented in the summary come from the source text (Attribution) (Figure 4) but it surprisingly does not follow the same behaviour as the metric trained to detect if the main idea is present (Main idea). Coincidentally, the main idea metric does not correlate with the error probability of the classification tasks. This may be either an artefact of the training of SEAHORSE benchmark, a limitation of our metric, or could indicate that answering the question "Is the content of the summary fully attributable to the content of the source text" is more relevant for downstream tasks than "Is the main idea of the source text present in the summary".
Independence of the metrics. While it seems that Attribution and the MI follow a similar trajectory in Figure 4, we found that the MI is not correlated with the Attribution metric (Figure 5). This suggests that these two metrics are independent and could be used in conjunction to evaluate summarizers. When it comes to more standard metrics, we find two clusters of metrics that seem to be relatively independent of each other: one represented by BERTScore and comprising the MI and the other represented by BARTScore which includes Attribution.
@@ 8 Summary and Discussion
We have demonstrated that the mutual information between source texts and summaries serves as a universal and task-agnostic measure of how well a summarizer preserves useful information for decision-making. A high mutual information indicates that a task performed on the summaries is likely to produce the same outcome as on the source texts. COSMIC is therefore theoretically grounded in a reasonable task-oriented scenario. Our ability to estimate this mutual information practically and its correlation with downstream task performance further underscores its significance.
Our proposed method extends beyond summarization systems and could also contribute to the broader field of multi-modal generation evaluation Kim et al. (2022). However, we acknowledge that there is still much to explore in multi-modal generation evaluation.
@@ Appendix A Proof of the Upper and Lower Bounds
A.1 Proof of the upper bound on the error probability
@@ Appendix B Rate-Distortion Bound for Classification Tasks
Note that the expected distortion equals the error probability, i.e.,
@@ Appendix C Model Specifications
sshleifer/distilbart-xsum-12-1 cnn_dailymail 221 M 0.04 -0.04 -3.16 0.21 0.41 0.87 0.83 0.22 0.67 49.70 -10.20 15.21
@@
sshleifer/distilbart-cnn-12-3 cnn_dailymail 255 M 0.18 0.24 -1.36 0.48 0.82 0.87 0.93 0.54 0.60 55.21 -15.71 -13.15
@@
xsum 255 M 0.31 0.32 -1.58 0.49 0.74 0.82 0.94 0.47 0.64 51.50 -17.45 -11.65
sshleifer/distilbart-cnn-12-6 xsum 305 M 0.31 0.33 -1.50 0.54 0.76 0.91 0.96 0.51 0.71 51.65 -17.58 -12.11
@@
cnn_dailymail 305 M 0.18 0.25 -1.31 0.52 0.84 0.91 0.93 0.58 0.62 55.51 -16.00 -13.79
sshleifer/distill-pegasus-cnn-16-4 cnn_dailymail 369 M 0.13 0.19 -1.40 0.48 0.83 0.81 0.92 0.54 0.64 54.69 -15.19 -8.70
@@
xsum 369 M 0.23 0.28 -1.49 0.47 0.81 0.84 0.94 0.49 0.70 51.32 -17.29 -6.70
sshleifer/distill-pegasus-xsum-16-4 cnn_dailymail 369 M 0.04 -0.04 -2.92 0.27 0.50 0.96 0.95 0.32 0.85 46.35 -6.84 13.27
xsum 369 M 0.08 0.05 -2.88 0.44 0.45 0.97 0.97 0.36 0.91 46.65 -12.57 16.31
facebook/bart-large-cnn xsum 406 M 0.32 0.34 -1.31 0.52 0.81 0.96 0.97 0.53 0.74 52.36 -18.33 -13.06
@@
cnn_dailymail 406 M 0.18 0.25 -1.19 0.49 0.86 0.95 0.94 0.59 0.67 55.54 -16.05 -13.55
google/pegasus-multi_news xsum 570 M 0.27 0.20 -2.58 0.53 0.49 0.94 0.85 0.39 0.72 49.01 -15.00 -10.98
multi_news 570 M 0.13 0.09 -2.24 0.40 0.70 0.91 0.79 0.46 0.63 47.44 -13.40 -9.99
cnn_dailymail 570 M 0.16 0.16 -2.26 0.47 0.63 0.94 0.83 0.43 0.65 53.89 -14.39 -13.12
google/pegasus-large xsum 570 M 0.27 0.31 -1.19 0.24 0.88 0.86 0.89 0.37 0.70 47.92 -13.89 -4.18
multi_news 570 M 0.12 0.19 -1.35 0.17 0.82 0.72 0.77 0.39 0.47 48.58 -14.54 -9.38
cnn_dailymail 570 M 0.20 0.25 -1.21 0.26 0.87 0.66 0.84 0.43 0.51 53.73 -14.21 -9.28
@@
multi_news 570 M 0.06 -0.30 -3.20 0.06 0.34 0.29 0.43 0.13 0.37 39.79 -5.75 1.47
cnn_dailymail 570 M 0.10 -0.26 -3.30 0.08 0.35 0.26 0.45 0.12 0.30 44.37 -4.86 -0.28
mistralai/Mistral-7B-Instruct-v0.2 multi_news 7241 M 0.16 0.09 -2.84 0.37 0.64 0.88 0.80 0.41 0.68 45.30 -11.38 -9.15
cnn_dailymail 7241 M 0.25 0.25 -2.20 0.54 0.74 0.93 0.83 0.46 0.67 54.88 -16.04 -14.53
xsum 7241 M 0.42 0.35 -2.23 0.59 0.65 0.94 0.86 0.44 0.76 51.52 -18.28 -16.69

### v3

@@ 5.1 Downstream tasks
Classification tasks. We selected four different tasks (sentiment analysis, policy classification, Emotion classification and ChatGPT detector) and corresponding classifiers from the Huggingface Hub to run on the source texts and summaries and compared their outputs. We report the expected error rate(6), i.e., the classifier outputting a different label on the source text and the summary.
@@ 5.2 Correlation With Learnt Metrics
There are not many available metrics trained on human judgment for summarization. We chose to evaluate our metric on the SEAHORSE metrics Clark et al. (2023) as they are the most recent and provide interpretable metrics.
Seahorse metrics. These metrics — learnt from human judgement — assess summaries along $6$ axes: Main ideas; Attribution; Comprehensible; Grammar; Repetition and Conciseness. Main ideas and Attribution, respectively, measure if the main ideas of the source text are present in the summary and if the information in the summary does indeed come from the source text. Comprehensible and Grammar measure if the summary is fluent and grammatically correct, while Repetition and conciseness measure if there are NO repetitions and the summary is concise. The $\Pr(\text{Yes})$ metric corresponds to the average probability over the dataset that the SEAHORSE model predicts the answer Yes to the corresponding question. While we would not expect our MI metric to correlate with the Grammar or Comprehensible scores, it should strongly correlate with the Main Ideas and Attribution scores as these are proxies to the information contained in the summary.
@@ 6 Experimental Results
Table 1: Common quality estimation metrics correlation with the performance on the downstream classification tasks. Where Sent. analysis stands for sentiment analysis, GPT det. for GPT detector, Topic. for topic classification, Policy for policy classification, Emotion for emotion classification and Emb. for paraphrase embedding.
@@
$I(S;T)$ 0.63 0.59 0.58 0.56 0.81
Bas. BERTScore Precision 0.53 0.68 0.47 0.46 0.70
BERTScore Recall 0.59 0.66 0.74 0.54 0.42
BLANC 0.59 0.59 0.66 0.60 0.38
SMART1 0.55 0.63 0.47 0.47 0.28
SMART2 0.55 0.63 0.47 0.47 0.28
SMARTL 0.55 0.63 0.47 0.47 0.28
BARTScore 0.51 0.73 0.42 0.48 0.62
BERTScore 0.64 0.75 0.64 0.58 0.61
ROUGE-L 0.56 0.55 0.54 0.47 0.29
SH. Attribution 0.49 0.71 0.37 0.49 0.62
Main ideas 0.33 0.37 0.60 0.38 0.47
@@
Correlation with downstream tasks performance. In Table 1, we report the correlation between the different metrics and the expected error rate for different classification tasks and with the dot product for the Paraphrase embedding task. The MI is competitive with both the common quality estimation metrics such as BERTScore and BARTScore. In addition, we report the correlation of the metrics trained on human judgement from the SEAHORSE benchmark to predict whether the main idea of the text is present in the summary and if all elements of the summary are attributable to the source text.
@@
Figure 2: Spearman correlation with human judgment estimated by the SEAHORSE metrics. As one would expect the MI does correlate with Attribution and Main ideas but not with comprehensible, grammar or repetition.
@@
Correlation with human-judgement-based metrics. As the SEAHORSE metrics are trained to mimic human judgment, we can use them to assess the behavior of the MI. We found consistent and expected correlations with the different SEAHORSE metrics (Figure 2). The mutual information correlates well with Attribution and Main Ideas but not with Comprehensible, Grammar and Repetition. This is not surprising as one would expect the mutual information to capture the amount of information in the summary that is attributable to the source text and not the grammatical correctness or fluency of the summary. The high correlation with Conciseness is a rather surprising result. We believe this is because the conciseness correlates with the strength of the summarizer, which correlates with the MI between the source texts and the summaries. The stronger the language model, the better the source texts’ encoding as a summary. It is also plausible that the learned metrics are flawed in some ways, hindering the results.
@@
We observe that the MI correlates more consistently with these Main idea, Attribution and Conciseness scores than the common quality assessment metrics, e.g. BERTScore and BARTScore (see Figure 2). It suggests that in addition to being a good predictor of the performance of the downstream tasks, MI is also a better predictor of the human judgment of the quality of the summaries. Notably, the MI, a theoretical quantity derived from Shannon’s MI, reproduces human judgment expectations without training on human judgment data.
@@
Comparing summarizers with COSMIC. In Figure 3, we compare the MI of different models and report their size. We observe that OOD models — trained on Arxiv or medical data — perform very poorly, whereas IN distribution models such as BART display significantly higher MI. Interestingly enough, the size of the model does not seem to be a good predictor of MI. In Appendix F we explore further use of the mutual information to compare the informativeness of different summarizers between themselves to construct a hierarchy based on their mutual informativeness.
@@
Figure 3: Comparison of models of different sizes and strengths, finetuned on different datasets regarding the MI between their summaries and the source text on different evaluation datasets. The dotted lines highlight the highest MI for each dataset reached by a model, and the grey area represents the number of parameters of the model
Metric Coher. Cons. Flu. Rel.
Reference Dependent
ROUGE-1 .35 .55 .53 .58
ROUGE-2 .23 .60 49 .43
ROUGE-L .12 .12 .26 .35
BLEU .22 .05 .33 .38
CHRF .35 .63 .56 .55
BERTScore .33 -.03 .14 .20
MoverScore .23 -.05 .26 .35
BLEURT .53 .20 .41 47
SMS .27 60 .36 .40
SMART-1 .43 67 .64 .67
SMART-2 .42 .75 .63 .58
SMART-L .57 .57 .61 .73
Reference Free
PRISM .23 .60 .36 .37
T5-ANLI .25 .58 .54 .52
BARTScore .35 .62 .49 45
BARTScore+CNN .55 .32 .59 .58
$Q^{2}$25 .75 .58 .45
RISE ${}_{\text{extMulti-News }}$53 .73 .71 .70
RISE ${}_{\text{SamSUM }}$53 .70 .68 .70
$\operatorname{RISE}_{CNN}$53 .73 .75 .70
Ours
$I(\textbf{T};\textbf{S})$23 .53 .47 .54
Table 2: Comparison of our method against many baselines on the SummEval Human evaluation dataset. We report the system-level Kendall’s Tau correlation with human judgments.
7 Discussion of Quantitative Results
@@
Our findings reveal that various conventional metrics do not consistently align with the effectiveness on downstream tasks. As illustrated in Figure 4, different metrics exhibit distinct behaviors and correlations with other metrics and the performance of downstream tasks.
@@
BERTScore. BERTScore displays good correlation with the task-preserving capabilities of the summaries but have poor correlations with human judgements (real or estimated); whereas the mutual information is theoretically grounded and an all-around more consistent metric.
Figure 4: Correlation between the main metrics and different performance metrics for the different datasets. The MI (green) closely follows the behavior of the Attribution metric but is a better predictor of the performance of the downstream tasks (Sentiment analysis, GPT Detector, Topic classification, etc.). By contrast, ROUGE scores do not display consistent correlations. For metrics to be considered effective, they should consistently exhibit positive correlations with each other. In Section E.3 we report the aggregated numerical results.
@@
Behavior of MI. The MI displays very similar behavior to the metric trained to detect whether the ideas presented in the summary come from the source text (Attribution) (Figure 4), but it surprisingly does not follow the same behavior as the metric trained to detect if the main idea is present (Main idea). Coincidentally, the main idea metric does not correlate with the expected error rate of the classification tasks. This may be an artifact of the training of SEAHORSE benchmark, a limitation of our metric, or could indicate that answering the question "Is the content of the summary fully attributable to the content of the source text" is more relevant for downstream tasks than "Is the main idea of the source text present in the summary".
Independence of the metrics. While it seems that Attribution and the MI follow a similar trajectory in Figure 4, we found that the MI is not correlated with the Attribution metric (Figure 5). This suggests that these two metrics are independent and could be used in conjunction to evaluate summarizers. When it comes to more standard metrics, we find two clusters of metrics that seem to be relatively independent of each other: one represented by BERTScore and comprising the MI and the other represented by BARTScore, which includes Attribution.
Correlations with Human Judgement. While the MI correlates less on the SummEval judgment dataset than competitors, such as SMART or RISE, it correlates more with the downstream tasks performance (Table 1) while not relying on golden summaries for comparison (SMART) or on very large pre-trained models (RISE). As mentioned previously, the comparison with other metrics is not completely fair as the MI is evaluated on unannotated data while the other metrics are evaluated on the (limited) annotated data. Nonetheless, the MI displays on-par performance with the other metrics on the SummEval dataset and is a promising metric for summarizer evaluation overall.
@@ 8 Summary and Discussion
We introduced a task-oriented setting for summary evaluation in which we can derive a principled and clear notion of the quality of a summarizer: the expected risk of performing a task on a summary instead of the source text. We connected this risk theoretically to the mutual information of the source texts and generated summaries and we bounded the risk on both sides using the mutual information. Even if these bounds are task-agnostic and thus potentially loose, we demonstrated experimentally that the mutual information indeed correlated well with the risk. High mutual information indicates that a task performed in the summaries is likely to produce the same outcome as in the source texts. COSMIC is therefore theoretically grounded in a reasonable task-oriented scenario. Our ability to estimate this mutual information practically and its correlation with downstream task performance further underscores its significance. Our proposed method extends beyond summarization systems and could also contribute to the broader field of multi-modal generation evaluation Kim et al. (2022).
Acknowledgements
This work was granted access to the HPC resources of IDRIS under the allocation 2023-AD011013290R2 made by GENCI.
@@ Appendix A Proof of the Upper and Lower Bounds
A.1 Proof of the upper bound on the expected error rate
@@ Appendix B Rate-Distortion Bound for Classification Tasks
Note that the expected distortion equals the expected error rate, i.e.,
@@ Appendix C Model Specifications
All the evaluated models are listed with their characteristics in Table 3.
@@
sshleifer/distilbart-xsum-12-1 xsum 221 M 0.08 0.03 -3.27 0.29 0.35 0.86 0.90 0.22 0.80 46.52 -12.49 17.01
@@
sshleifer/distilbart-cnn-12-3 xsum 255 M 0.31 0.32 -1.58 0.49 0.74 0.82 0.94 0.47 0.64 51.50 -17.45 -11.65
cnn_dailymail 255 M 0.18 0.24 -1.36 0.48 0.82 0.87 0.93 0.54 0.60 55.21 -15.71 -13.15
@@
sshleifer/distilbart-cnn-12-6 cnn_dailymail 305 M 0.18 0.25 -1.31 0.52 0.84 0.91 0.93 0.58 0.62 55.51 -16.00 -13.79
@@
xsum 305 M 0.31 0.33 -1.50 0.54 0.76 0.91 0.96 0.51 0.71 51.65 -17.58 -12.11
sshleifer/distill-pegasus-xsum-16-4 xsum 369 M 0.08 0.05 -2.88 0.44 0.45 0.97 0.97 0.36 0.91 46.65 -12.57 16.31
cnn_dailymail 369 M 0.04 -0.04 -2.92 0.27 0.50 0.96 0.95 0.32 0.85 46.35 -6.84 13.27
sshleifer/distill-pegasus-cnn-16-4 xsum 369 M 0.23 0.28 -1.49 0.47 0.81 0.84 0.94 0.49 0.70 51.32 -17.29 -6.70
@@
cnn_dailymail 369 M 0.13 0.19 -1.40 0.48 0.83 0.81 0.92 0.54 0.64 54.69 -15.19 -8.70
facebook/bart-large-cnn cnn_dailymail 406 M 0.18 0.25 -1.19 0.49 0.86 0.95 0.94 0.59 0.67 55.54 -16.05 -13.55
@@
xsum 406 M 0.32 0.34 -1.31 0.52 0.81 0.96 0.97 0.53 0.74 52.36 -18.33 -13.06
google/pegasus-multi_news multi_news 570 M 0.06 0.02 -2.51 0.48 0.69 0.93 0.79 0.52 0.58 46.71 -12.68 -4.24
xsum 570 M 0.27 0.20 -2.58 0.53 0.49 0.94 0.85 0.39 0.72 49.01 -15.00 -10.98
@@
google/pegasus-large cnn_dailymail 570 M 0.20 0.25 -1.21 0.26 0.87 0.66 0.84 0.43 0.51 53.73 -14.21 -9.28
multi_news 570 M 0.07 0.12 -1.52 0.18 0.82 0.82 0.72 0.37 0.42 47.86 -13.81 -5.25
xsum 570 M 0.27 0.31 -1.19 0.24 0.88 0.86 0.89 0.37 0.70 47.92 -13.89 -4.18
google/pegasus-multi_news cnn_dailymail 570 M 0.16 0.16 -2.26 0.47 0.63 0.94 0.83 0.43 0.65 53.89 -14.39 -13.12
google/pegasus-arxiv cnn_dailymail 570 M 0.10 -0.26 -3.30 0.08 0.35 0.26 0.45 0.12 0.30 44.37 -4.86 -0.28
multi_news 570 M 0.05 -0.27 -3.59 0.06 0.32 0.36 0.46 0.12 0.39 39.87 -5.82 1.21